APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Saturday January 28, 2006 (MJD 53764) ---=== OBSERVING TEAM ===--- Night: John Barentine Mike Harvanek Support: Eric Neilsen (phone) ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- A mixed night of mostly spectroscopy and a less-than-productive stint at taking the imager out for a spin. We finished four plates (2399, 2404, 2370, 2141), obtained partical signal on one plate (2374) and imaged SEGUE stripe 1540 from +147 <= Mu <= +162. The imaging run was in marginal conditions and may not be acceptable. A serious failure of the goDrift command in IOP caused us to lose about 45 minutes of rapidly-vanishing condtions for good imaging in the middle of the night. PR 6868 against IOP v4_12_0 was filed and the workaround is in the Problems section of this log. DIMM lasted for about a total of one hour tonight; we shut it down after finishing the first plate of the night when it stopped reporting seeing values. ---=== OBSERVING LOG ===--- Afternoon --------- Spectro focus information is in the Problems section. No trouble was encountered during checkout. Night ----- Mostly clear and cold at sunset, airTemp=-3 C, winds moderate out of the NNW near 20MPH and humidity low. Open at 00:40Z with a slight delay thanks to your humble correspondent, who neglected to ensure the building pins were removed before hitting the big "go" button. Cleared the overload, back up slightly to un-wedge the pins to where we could pull them. The building came off just fine after that. Mike had the "can't take UDP port 30410" problem tonight when he tried to take the instruments from me before we got on the sky. I had to log completely out of my SOP session in order to allow him to take the instruments. The cloud camera suddenly died around 01:25Z. We did a scanstop/scanstart and it was back running again. 01:30Z Cart 1 Plate 2399 (tile 8149, seguefaint134). GSOGTF in twilight. Seeing was good to start, often better than one arcsecond, and our first exposure for 1200 s gave a low (S/N)^2 of 8.1. The efficiency was a little lower starting with the next exposure, which we upped from 20 min to 25 min. SoS reported whopping fibers #40 and #23, both of which are bright (mag 14-15) standards. We continued upping the exposure times to 1800 s beginning with the third exposure, and did the SoS-recommended postCalib after the fourth exposure. The SoS results from the postCalib suggested our focus was drifting, so we decided to simply mechPiston sp2 on mounting the next cartridge without doing a formal goSpecFocus. This trick worked last night for sp1. Our strategy of extending the exposure times also worked, allowing us to finish this plate with a low total (S/N)^2 of 51.0 in r1. Approaching clouds from the west on satellite during this plate convinced us that the likelihood of the sky remaining photomteric for two hours after finishing Plate 2399 were slim to none. We continued with spectroscopy. 04:19Z Cart 6 Plate 2404 (tile 8154, seguefaint134) No GSOGTF but found an fk5InFiber number 10. After setting up on the field we did a "mechPiston -3000" on sp2 (final motor numbers are in the Problems section). That seemed to fix the problem on sp2 but we had a yellow WSIGMA result on the arc taken during the preCalib on this plate. We aborted the first science exposure a few minutes in to fix the focus on sp1 rather than lose the exposure were it go to out of spec during integration. So we issued a "mechPiston -3000" on sp1 after aborting the science exposure, took another preCalib to check the results, and started the science exposure over again. Note that neither the ambient airTemp outside nor the telescope temp had changed appreciably since we checked spectro focus in the afternoon. SoS warned about whopping fibers 38, 65, 160 and 603; all are bright standard stars > mag 15. Efficiency was good on this plate and we backed the fourth exposure off from 30 min to 20 min. Finished the plate with a low (S/N)^2 of 53.2, again in r1. 07:05Z Cart 7 Plate 2370 (tile 1696, chunk131) GSOGTF with an additional offset of A=-5 beyond the sticky offsets from the previous plate. Clouds started to come in from the SW during the setup on this plate. Done in 3 x 900 s exposures with a low (S/N)^2 of 15.8 (that pesky r1...) Seeing an opportunity to image, we took it. We appeared to be in a ribbon of clear sky between cloud bands separated (we believed) by about 2 to 2.5 hours. Switched at 08:15Z but then on setup of the imager we immediately had a problem issuing the goDrift command: im> goDrift Flavor of this run (science, bias, calibration, engineering, ignore, help) []: science Stripe, Strip (e.g, 22,N or SS,N) [17,N]: Please provide a one line comment [] go Error: called "eqToStripe" with too many arguments (A full traceback of the error is in the Problems section along with the solution.) Called Eric, who walked us through diagnostics and then a fix. This involved two aborted drifts (Run 5997 for one frame and Run 5997 for 3 frames). 09:10Z Run 5999 (SEGUE stripe 1540, startMu=+147) Skippy happy from the start, as was focus. Seeing was a little shaky at the beginning but trending down to within spec by about frame 20. Ramp began at Frame 22, Mu=+149.5, and science at Mu=+153. We continued imaging but we had both thin clouds moving in and seeing that was never reliably "good" according to seeingRun. We ended the drift at Frame 106 (Mu=+162). 10:07Z Run 6000 (On-telescope bias) Switched back to spectroscopy at 10:30Z. 10:40Z Cart 4 Plate 2374 (Tile 1700, chunk131) GSOGTF. Clouds and an early expiry of this plate made us pull it after one exposure of 1200 s. 11:35Z Cart 9 Plate 2141 (Tile 1547, chunk109). No GSOGTF but found an fk5InFiber number 10 with an additional offset of A=-4, D=+3. We did two exposures of 1200 and 600 s, bringing us up to about 1/2 hour before twilight. Without enough time to finish this plate, mount another cartridge, setup on the field and get in a single exposure before pushing too far into twilight. We decided to take an extra 600 s exposure to get additional signal on this plate, then call it a night. Finished with more than enough signal on this plate. Closed up for the night at 13:00Z. endNight rejoiced ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 5996 23:36Z 23:51Z 100 O -168.62 -164.94 39 ignore 5997 08:34Z 08:35Z 100 O -33.75 -33.50 16 ignore 5998 08:57Z 09:01Z 17 N -68.68 -67.82 19 ignore 5999 09:08Z 10:08Z 1540 S -37.71 -22.60 114 science 6000 10:12Z 10:26Z 100 O -9.13 -5.62 38 bias ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- 5998 8 4 0 0.000 0.000 0.000 0.000 -0.00000 0 0 5999 11 4 11 -9.100 1.132 9.500 1.980 0.05472 178 32 5999 29 4 87 6.800 0.231 1.800 0.183 -0.02069 180 35 5999 47 4 75 6.300 0.190 1.600 0.131 -0.01816 181 37 5999 63 4 66 6.000 0.189 1.400 0.138 -0.01574 182 39 ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ 5999 22 1 34 1 39 -1.362 0.358 2.909 0.251 -0.00978 5999 22 1 34 1 39 -1.091 0.178 2.922 0.258 -0.00982 5999 22 1 34 1 26 -0.710 0.261 2.803 0.289 -0.00942 5999 22 1 34 1 30 -0.352 0.164 3.076 0.197 -0.01034 5999 22 1 34 1 28 -0.038 0.349 3.092 0.309 -0.01040 5999 22 1 34 1 39 -1.362 0.358 2.909 0.251 -0.00978 5999 22 1 34 1 39 -1.091 0.178 2.922 0.258 -0.00982 5999 22 1 34 1 26 -0.710 0.261 2.803 0.289 -0.00942 5999 22 1 34 1 30 -0.352 0.164 3.076 0.197 -0.01034 5999 22 1 34 1 28 -0.038 0.349 3.092 0.309 -0.01040 5999 38 1 36 359 29 -1.974 0.335 2.126 0.310 -0.00715 5999 38 1 36 359 21 -1.719 0.207 2.370 0.220 -0.00797 5999 38 1 36 359 25 -1.386 0.202 2.428 0.177 -0.00816 5999 38 1 36 359 27 -1.013 0.248 2.597 0.279 -0.00873 5999 38 1 36 359 23 -0.555 0.412 2.653 0.235 -0.00892 5999 72 1 41 356 16 -1.150 0.429 1.944 0.583 -0.00654 5999 72 1 41 356 25 -1.068 0.348 2.031 0.487 -0.00683 5999 72 1 41 356 18 -0.572 0.345 2.523 0.486 -0.00848 5999 72 1 41 356 26 -0.211 0.356 2.562 0.417 -0.00861 5999 72 1 41 356 14 0.089 0.283 2.863 0.194 -0.00963 5999 90 1 43 354 22 -0.747 0.238 2.027 0.249 -0.00681 5999 90 1 43 354 19 -0.365 0.333 2.078 0.216 -0.00699 5999 90 1 43 354 15 -0.241 0.253 2.283 0.211 -0.00768 5999 90 1 43 354 12 0.212 0.313 2.307 0.220 -0.00776 5999 90 1 43 354 19 0.784 0.268 2.357 0.281 -0.00792 ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): yes QA Procedures Done (y/n): yes UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 36823, plate -9999 ------- 23:21 36823 0.0 bias 23:25 36824 0.0 bias 23:28 36825 0.0 bias ----- sequence 36826, plate 0 ------- 23:31 36826 10.0 flat 23:33 36827 2.0 arc N.B.: The "Plate 0" was the result of not having loaded the cartridge before taking a postCalib, and the data were not reduced in SoS. The exposures were set to "test". ----- sequence 36828, plate 1987 ------- 23:56 36828 10.0 flat 23:58 36829 2.0 arc ----- sequence 36830, plate -9999 ------- 01:16 36830 0.0 bias ----- sequence 36831, plate 2399 ------- 77.1 51.0 93.7 61.5 DONE 01:36 36831 10.0 flat 01:38 36832 2.0 arc 02:05 36833 1200.1 target 02:33 36834 1500.1 target 03:00 36835 1500.1 target 03:33 36836 1800.1 target 03:36 36837 10.0 flat 03:38 36838 2.0 arc 04:12 36839 1800.1 target ----- sequence 36840, plate 2404 ------- 62.2 53.2 87.5 84.4 DONE 04:34 36840 10.0 flat 04:36 36841 2.0 arc 04:50 36842 10.0 flat 04:52 36843 2.0 arc 05:26 36844 1800.1 target 05:59 36845 1800.1 target 06:31 36846 1800.1 target 06:54 36847 1200.1 target 06:57 36848 10.0 flat 06:59 36849 2.0 arc ----- sequence 36850, plate 2370 ------- 23.2 15.8 22.1 20.4 DONE 07:12 36850 10.0 flat 07:14 36851 2.0 arc 07:34 36852 900.1 target 07:52 36853 900.1 target 08:09 36854 900.1 target ----- sequence 36855, plate 2374 ------- 5.0 3.7 5.2 4.0 10:55 36855 10.0 flat 10:58 36856 2.0 arc 11:26 36857 1200.1 target ----- sequence 36858, plate 2141 ------- 31.6 27.8 32.6 26.1 WAY DONE 11:39 36858 10.0 flat 11:42 36859 2.0 arc 12:06 36860 1200.1 target 12:18 36861 600.1 target 12:41 36862 1200.1 target 12:54 36863 600.1 target ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 02:11Z 1 2399 -68.47 -0.0019 73.41 0.0012 119.83 -0.0070 1.000140 05:05Z 6 2404 135.53 0.0018 51.69 0.0002 -77.13 -0.0223 1.000140 07:19Z 7 2370 46.10 0.0001 69.94 0.0020 219.48 0.0037 1.000050 11:07Z 4 2374 283.94 -0.0001 56.54 0.0006 119.72 -0.0075 1.000050 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL Stays: JL ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 02:12Z 1 2399 1.00014 -1383 157 175 -69 73.4 -3.0 16 316 -- 1.2 05:05Z 6 2404 1.00014 -1383 156 125 136 51.7 -2.3 17 300 -- 1.4 07:20Z 7 2370 1.00005 -494 798 75 46 70.0 -1.8 14 292 -- 1.3 11:07Z 4 2374 1.00005 -494 803 50 284 56.5 -1.4 22 300 -- 1.6 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 23:02Z 27 5 19 347 (NNW) 426 - MstlyClr 23:32Z 26 6 6 0 (N) 406 - MstlyClr 00:02Z 26 5 8 326 (NW) 418 - MstlyClr 00:32Z 25 5 13 329 (NNW) 386 - MstlyClr 01:03Z 26 4 18 321 (NW) 406 - Clear 01:33Z 26 4 18 308 (NW) 388 - Clear 02:03Z 26 5 20 312 (NW) 340 - Clear 02:34Z 26 6 15 302 (WNW) 304 - Clear 03:06Z 26 6 11 315 (NW) 365 - Clear 03:38Z 26 6 15 286 (WNW) 417 - Clear 04:09Z 27 5 14 301 (WNW) 595 - Clear 04:39Z 27 6 9 285 (WNW) 706 - Clear 05:10Z 28 5 12 292 (WNW) 774 - Clear 05:40Z 28 6 17 306 (NW) 758 - Clear 06:12Z 29 5 18 325 (NW) 640 - Clear 06:43Z 29 4 25 313 (NW) 525 - MstlyClr 07:13Z 29 4 14 315 (NW) 423 - MstlyClr 07:44Z 27 4 17 311 (NW) 466 - MstlyClr 08:14Z 27 5 8 319 (NW) 530 - MstlyClr 08:44Z 29 8 18 311 (NW) 238 - MstlyClr 09:14Z 28 7 16 326 (NW) 244 - PtlyCldy 09:45Z 28 8 20 316 (NW) 156 - PtlyCldy 10:15Z 29 8 24 305 (NW) 245 - PtlyCldy 10:47Z 29 7 25 310 (NW) 311 - PtlyCldy 11:17Z 29 7 22 294 (WNW) 291 - MstlyCldy 11:48Z 28 7 18 296 (WNW) 350 - MstlyCldy 12:19Z 29 8 22 298 (WNW) 382 - PtlyCldy 12:50Z 29 7 24 292 (WNW) 431 - PtlyCldy 13:26Z 29 6 17 285 (WNW) 407 - ---=== TELESCOPE STATUS ===--- 00:15Z Doors cracked, fans on, louvers open 00:40Z Enclosure off 13:00Z Enclosure on, doors and louvers closed, fans off Telescope parked at 121,30 with Cart 8 (Plate 2417) mounted Counterweights at 245 Spectro autofill reconnected No interlocks bypassed Dewar weights at 13:20Z Imager: 214lbs 22 psi Spectro: 230lbs 16 psi (force-filling) ---=== SOFTWARE USED ===--- IOP/SOP: v4_12_0 Watcher: v2_31_0 MCP: v5_26_2 TPM: tpm_v3_3_0 AstroDa: v15_10 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v2_09 SoS: v4_10_7 hoggPT: v1_6_9 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.8430 -5.8440 Axial B -5.6830 -5.6770 Axial C 0.9330 0.9420 Trans D -9.9910 -9.9940 Lateral E 10.3759 10.3759 Lateral F 11.9634 11.9380 GALILS Commanded: 4700. -2450. 3000. -7800. 3300. 3000. Actual: 4711. -2457. 2997. -7825. 3296. 3024. SETMIR VALUES PriDesOrient: 0.00 -1.68 7.05 385.25 92.02 PriOrient: 0.00 -1.52 7.02 384.99 93.81 SECONDARY: ---------- Focus: 0.00 Air Temp.: -2.9 Alt.: 30.363842 MIGS TONIGHT NOMINAL Axial A 1.5410 1.5420 Axial B 1.0600 1.0470 Axial C 1.1750 1.1650 Trans D -0.1200 -0.1240 GALILS Commanded: 1593356. 1603460. 1554055. -6800. -5900. Actual: 1593447. 1603642. 1554316. -6820. -5882. SETMIR VALUES SecDesOrient: 1257.00 15.87 0.00 -50.00 136.76 SecOrient: 1256.86 15.91 -0.04 -49.58 136.71 ---=== PROBLEMS IN DETAIL ===--- Spectro Focus ------------- airTemp outside at time of focus check (23:30Z) is about the same as the telescope temp, -2.8 C. tpm_TM_PMSSAvg -2.78286 tpm_TM_TelairAvg.VALA -0.6925 Overnight low is predicted to be approximately -4 deg C. With this in mind, we took a postCalib to check that all the SoS results were reasonably within spec. I forgot to load the plate before taking the postCalib and got SoS warning about unable to reduce flat (need plug-map). I loaded the cartridge and tried an sos_redo, but got warnings like this: % Compiled module: REMOVE2REDO. REMOVE2REDO: Selecting MJD=53764 REMOVE2REDO: Trying to lock the logfile /data/spectro/spectrologs/53764/logfile-53764.fits REMOVE2REDO: Reading the logfile /data/spectro/spectrologs/53764/logfile-53764.fits MRDFITS: Binary table. 6 columns by 16 rows. % READFITS: FITS header has NAXIS or NAXISi = 0, no data array read MRDFITS: Binary table. 19 columns by 12 rows. MRDFITS: Null image, NAXIS=0 MRDFITS: Null image, NAXIS=0 MRDFITS: Null image, NAXIS=0 but the images themselves looked fine. We just took another postCalib rather than worrying about it. At time of focus check, motor numbers were: sp1 Coll_motor_A 4752 Coll_motor_B 4749 Coll_motor_C 4738 sp2 Coll_motor_A -5207 Coll_motor_B -8491 Coll_motor_C -7813 For the focus update made on sp2 during Cart 6 tonight.... Before: Coll_motor_A -5207 Coll_motor_B -8491 Coll_motor_C -7813 sp2> mechPiston -3000 After: Coll_motor_A -8213 Coll_motor_B -11493 Coll_motor_C -10824 ...and the subsequent focus update on sp1: Before: Coll_motor_A 4752 Coll_motor_B 4749 Coll_motor_C 4738 sp1> mechPiston -3000 After: Coll_motor_A 1751 Coll_motor_B 1753 Coll_motor_C 1738 goDrift Balks At Starting An Imaging Run: Description and Fix ------------------------------------------------------------- Tonight IOP failed to execute a requested goDrift command: im> goDrift Flavor of this run (science, bias, calibration, engineering, ignore, help) []: science Stripe, Strip (e.g, 22,N or SS,N) [17,N]: Please provide a one line comment [] go Error: called "eqToStripe" with too many arguments IOP error traceback: im> tb called "eqToStripe" with too many arguments while executing "eqToStripe $ra $dec -segue" invoked from within "lindex [eqToStripe $ra $dec -segue]..." invoked from within "keylset output stripe [lindex [eqToStripe $ra $dec -segue]..." invoked from within "if {[info exists segue]} { keylset output stripe [lindex [eqToStripe $ra $dec -segue] 0] } else { keylset output stripe [lindex [eqToStripe $ra ..." (procedure "getStripeFromTCC" line 23) invoked from within "getStripeFromTCC -segue" invoked from within "set segueReferenceList [getStripeFromTCC -segue]..." (procedure "checkStripeAgainstTCC" line 21) invoked from within "checkStripeAgainstTCC" invoked from within "if {![checkStripeAgainstTCC]} { puts "Input stripe and strip do not match actual telescope pointing." puts -nonewline "Continue anyway? " ..." invoked from within "if {[cequal $scanType(flavor) science]} { if {![checkStripeAgainstTCC]} { puts "Input stripe and strip do not match actual telescope pointi ... " (procedure "goDrift" line 38) invoked from within "goDrift" How to solve it (filed with PR 6868): Before issuing a goDrift command, do a source $env(IOP_DIR)/etc/checkLogs.tcl in the same iop window that goDrift will be issued. All goDrifts from this window will then work. A goDrift from another iop window will not work unless you give the source command in that window as well. Note that this problem does not affect goDrift -bias.